# A Transferable Physics-Informed Framework for Battery Degradation Diagnosis, Knee-Onset Detection and Knee Prediction

### v1

@@ 3.3 A case study: advanced battery management system functions in a performance digital twin
The histogram-based feature engineering method, hybrid physics-informed model, and the proposed fine-tuning strategy are key enablers for a concept of battery performance digital twin (PDT) or cloud battery management system (BMS) [55]. They can enable advanced BMS functionalities, such as online degradation diagnosis and prognosis, aging-aware battery classification, and second-life repurposing, as illustrated in Fig. 10. Time-series voltage and current data are commonly measured for all cells connected in series inside a battery pack. The onboard BMS can first aggregate these time-series voltage and current data as histograms and then communicate to the PDT on request or at a very slow sampling rate via the Internet-of-Things (IoT) gateway [55]. The global battery PDT, or the hybrid physics-informed model in this work, first uses a small amount of labeled data to create a local PDT in a target scenario. The local PDT is then used to estimate degradation modes and detect the degradation phase for each cell in a battery pack. Based on cell-level estimated degradation modes and degradation phase, the cell-to-cell heterogeneity inside a battery pack can be determined for aging-aware battery classification later on. For example, the sample cell [E4C] in Fig. 8 was detected to be in Phase 3. The pack or module where the cell is located may be either repurposed to second-life applications in which the knee occurrence can be stopped or recycled. The specific second-life repurposing, however, also requires additional information, such as battery energy and power capabilities, technical requirements of second-life applications, residual value estimation, etc [56].
Figure 10: The physical battery and its performance digital twin in the cloud.
@@ 4 Conclusions
To alleviate the technical, economic, and safety concerns arising from capacity knee occurrence during the service life of a battery, a transferable physics-informed framework that consists of a histogram-based feature engineering method, a hybrid physics-informed model, and a fine-tuning strategy was proposed for online battery degradation diagnosis and knee-onset detection. Specifically, pre-trained hybrid physics-informed models were first developed using 1D or 2D histogram data and their battery degradation mode estimation and phase detection performance were evaluated in the source scenarios using a scenario-aware pipeline. The pre-trained hybrid models were then fine-tuned using different amounts of labeled data, and deployed in the target scenario. Among the 5 histogram-based feature sets investigated, it was demonstrated that the current-voltage feature set was the best for battery degradation mode estimation in both source and target scenarios, here and possibly in field applications as well. The fine-tuning strategy was proven to be effective in improving not only battery degradation mode estimation but also degradation phase detection using 3 estimated degradation modes in the target scenario. With degradation phases detected with high accuracy, online prediction of battery capacity knee points can also be achieved by leveraging the strong linear correlation identified between knee-onset and knee points. Lastly, it has been found that using one cell with knee occurrence in the target scenario may be enough to achieve a satisfactory model accuracy.
As key enablers for the concept of battery performance digital twin (DPT) in the cloud, our proposed framework can enable advanced BMS functions, such as online degradation diagnosis and prognosis, aging-aware battery classification, and second-life repurposing. As a result, the overall value of electric vehicle batteries can be maximized before recycling. In terms of future work, it would be interesting to quantify the aleatoric uncertainty arising from the noisy data and the epistemic uncertainty from the model structure using the Bayesian approach.

### v2

@@ 3.2 Model evaluation in the target scenario
In Ref. [56], a physics-informed neural network (PINN) is proposed for battery state-of-health (SoH) estimation, and a fine-tuning strategy is proposed to transfer the PINN model for SoH estimation of batteries with different chemistries and charge/discharge protocols. The key differences between Ref. [56] and our work are specified as follows: 1) Problems & Models: The PINN is used to capture battery capacity fade dynamics and estimate battery state-of-health (SoH) in Ref. [56], while in our work, the PINN is used to capture battery degradation mode dynamics, i.e., loss of lithium inventory (LLI) and loss of active material (LAM), and its estimated degradation modes are then taken as inputs to an XGBoost model to perform battery degradation phase detection and online battery classification; 2) Input features & Data: The cycle-based features, i.e., mean, standard deviation, kurtosis, skewness, charging time, accumulated charge, curve slope, and curve entropy, extracted from charge data before the battery is fully charged, are used as the inputs to the PINN for SoH estimation in Ref. [56], while the histogram-based features, i.e., time spent within specific voltage and/or current ranges, extracted from full charge-discharge data regardless of battery usage, are used as the inputs to the PINN for degradation mode estimation in our work; 3) Outputs: The output of the PINN model is battery capacity in Ref. [56], while the outputs of our model (PINN+XGBoost) are three degradation modes, degradation phase. With accurately detected degradation phase, the knee-onset point can also be detected online as the transition point from Phase 1 to Phase 2. The advantages of our method are as follows: 1) The capacity fade is attributed to the growth of degradation modes (LLI and LAM). Our method is capable of estimating these degradation modes, which provides diagnostic information for the occurrence of battery capacity knee. With these estimated degradation modes, degradation phases and knee-onset point can be detected in the target scenario (CC-CV charge/WLTP driving discharge profiles) with high accuracy; 2) Instead of extracting from either charge or discharge data, the 2D histogram features are extracted from full charge-discharge data regardless of battery usage, and are proven to be effective in estimating battery degradation modes, detecting degradation phase and knee-onset point.
Transformer-based models have shown excellent performance of estimating battery SoH in recent studies [57][58]. However, the proposed hybrid physics-informed model and Transformer-based models offer distinct advantages in battery SoH estimation and degradation diagnosis. The hybrid physics-informed model aims to embed battery degradation dynamics into the neural network (NN) architecture, leveraging the advantages of both, i.e., the flexibility of NNs and the interpretability of physics. In contrast, Transformer-based models excel in battery SoH estimation accuracy without physical insights. The choice between the hybrid physics-informed and Transformer-based models depends on the specific requirements in battery applications.
3.3 Computational requirements
The offline training and online deployment of the hybrid model (DeepHPM and XGBoost) are performed on a desktop (OS: Windows 10, Memory: 64 GB, CPU: Intel Core i7-13700, GPU: NVIDIA GeForce RTX 3070). At the offline training stage, the search time for the optimal hybrid model structure using one histogram-based feature set is approximately 10 hours. The optimal structure of the hybrid model can be found in Table 8 and Table 9B. The training time for the hybrid model using one histogram-based feature set is approximately 4 minutes. At the online deployment stage, the fine-tuning time to create one local model in the target scenario is approximately 30 seconds. The test (or inference) time per sample is in milliseconds, which is negligible compared to the battery degradation rate, the sampling rate, and the histogram aggregation rate.
@@ 3.4 A case study: advanced battery management system functions in a performance digital twin
The histogram-based feature engineering method, hybrid physics-informed model, and the proposed fine-tuning strategy are key enablers for a concept of battery performance digital twin (PDT), or cloud battery management system (BMS) [59]. They can enable advanced BMS functionalities, such as online degradation diagnosis and prognosis, aging-aware battery classification, and second-life repurposing. Time-series voltage and current data are commonly measured for all cells connected in series inside a battery pack. The onboard BMS can first aggregate these time-series voltage and current data as histograms and then communicate to the PDT on request or at a very slow sampling rate via the Internet-of-Things (IoT) gateway [59]. The global battery PDT, or the hybrid physics-informed model in this work, first uses a small amount of labeled data to create a local PDT in a target scenario. The local PDT is then used to estimate degradation modes and detect the degradation phase for each cell in a battery pack. Based on cell-level estimated degradation modes and degradation phase, the cell-to-cell heterogeneity inside a battery pack can be determined for aging-aware battery classification later on. For example, the sample cell [E4C] in Fig. 7 was detected to be in Phase 3. The pack or module where the cell is located may be either repurposed to second-life applications in which the knee occurrence can be stopped, or be recycled. The specific second-life repurposing, however, also requires additional information, such as battery energy and power capabilities, technical requirements of second-life applications, residual value estimation, etc [60].
@@ 4 Conclusions
To alleviate the technical, economic, and safety concerns arising from capacity knee occurrence during the service life of a battery, a transferable physics-informed framework that consists of a histogram-based feature engineering method, a hybrid physics-informed model, and a fine-tuning strategy was proposed for online battery degradation diagnosis and knee-onset detection. Specifically, pre-trained hybrid physics-informed models were first developed using 1D or 2D histogram-based feature sets and their battery degradation mode estimation and phase detection performance were evaluated in the source scenarios using a scenario-aware pipeline. The pre-trained hybrid models were then fine-tuned using different amounts of labeled data, and deployed in the target scenario. Among the five histogram-based feature sets investigated, it was demonstrated that the 2D histogram-based 17-feature set was the best for battery degradation mode estimation in both source and target scenarios, here and possibly in field applications as well. The fine-tuning strategy was proven to be effective in improving not only battery degradation mode estimation but also degradation phase detection using 3 estimated degradation modes in the target scenario. With degradation phases detected with high accuracy, online prediction of battery capacity knee points can also be achieved by leveraging the strong linear correlation identified between knee-onset and knee points. Lastly, it has been found that using one cell with knee occurrence in the target scenario may be enough to achieve a satisfactory model accuracy.
As key enablers for the concept of battery performance digital twin (DPT) in the cloud, the proposed framework can enable advanced BMS functions, such as online degradation diagnosis and prognosis, aging-aware battery classification, and second-life repurposing. As a result, the overall value of electric vehicle batteries can be maximized before recycling. In terms of future work, 1) it would be interesting to quantify the aleatoric uncertainty arising from the noisy data and the epistemic uncertainty from the model structure using the Bayesian approach; 2) we will further evaluate the performance of the proposed framework in real-world applications (e.g., electric vehicles and grid storage) once battery data in the field are available; 3) Considering abundant unlabeled data in the field that poses new challenges to the fine-tuning strategy. Active learning algorithms that iteratively expand the labeled data in the lab will be investigated; 4) For different types of batteries or accelerated aging mechanisms, we will conduct a systematic study to quantify the accuracy loss of the proposed fine-tuning strategy once degradation modes of other types of batteries or accelerated aging mechanisms become available.
@@ Acknowledgements
Appendix A Supplementary figures
Figure 8: Hybrid physics-informed model.
Appendix B Supplementary tables
Table 8: Optimal DeepHPM structures for battery degradation mode estimation in the source scenario.
Feature set Target variable LLI LAM_NE LAM_PE
1D voltage-based 3-feature set [2, 64] [6, 32] [4, 64]
1D voltage-based 5-feature set [2, 56] [2, 56] [6, 48]
1D current-based 3-feature set [4, 64] [2, 48] [8, 40]
1D current-based 5-feature set [4, 56] [2, 40] [6, 64]
2D current-voltage 17-feature set [2, 32] [2, 32] [4, 48]
$[x,y]$ denotes the number of hidden layers and neurons per layer.
Table 9: Optimal XGBoost structures for battery degradation phase detection in the source scenario.
Hyperparameter Value
Number of trees 100
Learning rate 0.0207
Maximum depth of a tree 8
Minimum sum of instance weight 7
Table 10: DeepHPM degradation mode estimation performance in the target scenario.
Feature set Model Pre-trained Fine-tuned with 1 cell without knee occurrence Fine-tuned with 1 cell with knee occurrence Fine-tuned with 2 cells 1D voltage-based, 3-feature set [0.0465,0.0413,0.0423] [0.0457,0.0457,0.0423] [0.0329,0.0324,0.0269] [0.0228,0.0247,0.0307] 1D voltage-based, 5-feature set [0.0474,0.0454,0.0413] [0.0465,0.0470,0.0426] [0.0357,0.0343,0.0269] [0.0218,0.0198,0.0261] 1D current-based, 3-feature set [0.0468,0.0377,0.0401] [0.0456,0.0451,0.0428] [0.0386,0.0359,0.0304] [0.0337,0.0327,0.0304] 1D current-based, 5-feature set [0.0442,0.0385,0.0425] [0.0459,0.0447,0.0428] [0.0362,0.0302,0.0300] [0.0256,0.0268,0.0276] 2D current-voltage, 17-feature set [0.0670,0.0595,0.0570] [0.0478,0.0482,0.0426] [0.0381,0.0404,0.0371] [0.0214,0.0210,0.0265] $[x,y,z]$ denotes standard deviation values for LLI, LAM_NE, and LAM_PE, respectively. The bold values denote the minimum standard deviation.
Table 11: Robustness analysis of DeepHPM degradation mode estimation in the target scenario.
Case Model Pre-trained Fine-tuned with 1 cell without knee occurrence Fine-tuned with 1 cell with knee occurrence Fine-tuned with 2 cells 2D current-voltage, 17-feature set [0.0764,0.0597,0.0590] [0.0524,0.0510,0.0477] [0.0373,0.0406,0.0443] [0.0222,0.0220,0.0281] Additive Gaussian noise, $\mathcal{N}(0,0.04\sigma^{2}_{x})$ [0.1061,0.0723,0.0672] [0.0534,0.0532,0.0476] [0.0538,0.0630,0.0481] [0.0269,0.0267,0.0297] Additive Gaussian noise, $\mathcal{N}(0,0.25\sigma^{2}_{x})$ [0.1101,0.0791,0.0682] [0.0568,0.0592,0.0487] [0.0677,0.0772,0.0598] [0.0395,0.0407,0.0358] Missing histogram data [0.0768,0.0589,0.0583] [0.0518,0.0531,0.0475] [0.0473,0.0299,0.0570] [0.0237,0.0245,0.0282] $[x,y,z]$ denotes RMSE values for LLI, LAM_NE, and LAM_PE, respectively. The bold values denote the minimum RMSE. $\sigma_{x}$ denotes the standard deviation per input feature estimated in the source scenario.
Table 12: Gaussian process regression degradation mode estimation performance in the source scenario.
Feature set Target variable LLI LAM_NE LAM_PE
1D voltage-based 3-feature set 0.0061 0.0119 0.0067
1D voltage-based 5-feature set 0.0046 0.0135 0.0075
1D current-based 3-feature set 0.0069 0.0093 0.0084
1D current-based 5-feature set 0.0078 0.0094 0.0086
2D current-voltage 17-feature set 0.0046 0.0138 0.0073
Bold values denote the minimum RMSE in each column.
Table 13: Gaussian process regression degradation mode estimation performance in the target scenario.
Feature set Model Pre-trained
1D voltage-based, 3-feature set [0.0359,0.0328,0.0347]
1D voltage-based, 5-feature set [0.0395,0.0380,0.0337]
1D current-based, 3-feature set [0.0407,0.0422,0.0441]
1D current-based, 5-feature set [0.0501,0.0525,0.0481]
2D current-voltage, 17-feature set [0.0368,0.0333,0.0377]
$[x,y,z]$ denotes RMSE values for LLI, LAM_NE, and LAM_PE, respectively. The bold values denote the minimum RMSE.